# Higher-Dimensional Moving Averages and Submanifold Genericity

### v1

@@ Abstract.
We generalize results of Bellow, Jones, and Rosenblatt on moving ergodic averages to measure-preserving actions of $\mathbb{Z}^{d}$ and $\mathbb{R}^{d}$ for $d\geq 1$. In particular, we give necessary and sufficient conditions for the pointwise convergence of certain sequences of functions defined by averaging over families of boxes in $\mathbb{Z}^{d}$ and $\mathbb{R}^{d}$. As an application of our characterization, we show that averages along dilates of "locally flat" submanifolds in $\mathbb{R}^{d}$ do not necessarily converge point-wise for bounded measurable functions. This is closely related to the concept of submanifold-genericity recently introduced in [BFK25].
@@ 1. Introduction
Let $T$ be a measure-preserving transformation defined on a non-atomic probability space $(X,\mu)$. Let $\mathbb{N}$ denote the set of strictly positive integers and let $\Omega$ be a sequence of integer pairs $\{(n_{k},l_{k})\}\subset\mathbb{N}\times\mathbb{N}$. For each $k\in\mathbb{N}$ consider the averaging operator
$A_{k}f:=\frac{1}{l_{k}}\sum_{j=0}^{l_{k}-1}f\circ T^{n_{k}+j},$
where $f\in\textup{L}^{1}(X)$, and let
$M_{\Omega}f:=\sup_{k}\frac{1}{l_{k}}\sum_{j=0}^{l_{k}-1}\left|f\circ T^{n_{k}+j}\right|.$
In [BJR90], Bellow, Jones, and Rosenblatt gave necessary and sufficient conditions on the sequence $\Omega$ for $M_{\Omega}$ to satisfy a maximal inequality. Let us briefly recall their characterization. For $\alpha>0$ define
$\Omega^{(\alpha)}:=\left\{(x,y)\in\mathbb{N}^{2}:|x-n_{k}|\leq\alpha(y-l_{k})\mbox{ for some }(n_{k},l_{k})\in\Omega\right\}$
and for $\lambda\in\mathbb{N}$ set
$\Omega^{(\alpha)}(\lambda):=\left\{x\in\mathbb{N}:(x,\lambda)\in\Omega^{(\alpha)}\right\}.$
Then, according to [BJR90, Theorem 1], $M_{\Omega}$ is of weak type $(1,1)$ and of strong type $(p,p)$ for any $p>1$ if and only if the following condition on the sequence $\Omega$ is satisfied
($C_{1}$) $\exists\,A,\alpha>0:\quad\#\Omega^{(\alpha)}(\lambda)\leq A\lambda\quad\mbox{for all }\lambda\in\mathbb{N}.$
Condition ($C_{1}$) asserts in other words that the horizontal cross-section of all the cones of aperture $\alpha$ above points in the sequence $\Omega$ at fixed height must grow linearly as the height increases. For example, the sequence $(k,rk)$ for any $r\in\mathbb{N}$ has this property, while the sequence $(k,\sqrt{k})$ does not (see also picture below).
$\boldsymbol{0}$$l$$n$$(n_{1},l_{1})$$(n_{2},l_{2})$$(n_{3},l_{3})$$\lambda$$\Omega^{(1)}(\lambda)$
Figure 1. Set $\Omega^{(1)}(\lambda)$ for some sequence $(n_{k},l_{k})$.
It is a well-known fact that, when the operator $M_{\Omega}$ is of weak type $(1,1)$, the sequence of functions $A_{k}f$ converges point-wise for any $f\in\textup{L}^{1}(X)$. [BJR90, Theorem 4] shows additionally that this is not the case if Condition ($C_{1}$) fails.
In this paper, we generalize [BJR90, Theorem 1, Part a)] and [BJR90, Theorem 4] to measurable actions of $\mathbb{Z}^{d}$ and $\mathbb{R}^{d}$ for $d\geq 1$. Based on these extensions, we study the property of submanifold genericity, recently introduced in [BFK25]. We now proceed to lay out our results in greater detail.
1.1. The Discrete Case
Let $(X,\mu)$ be a non-atomic probability space and let $T_{1},\dotsc,\\ T_{d}$ be commuting measure-preserving transformations on the space $X$ (which is our standing assumption throughout this section). Given sequences $\{\boldsymbol{n}_{k}\}\subset\mathbb{Z}^{d}$ and $\{\boldsymbol{l}_{k}\}\subset\mathbb{N}^{d}$ ($k\in\mathbb{N}$), put
@@ 1.1. The Discrete Case
and let :=$\mathcal{B}:=\{B_{k}\}$. To any box $B_{k}\in\mathcal{B}$ associate an averaging operator $A_{k}$ defined by
@@
Our first result shows, roughly speaking, that, if Condition ($C_{1}$) holds in each co-ordinate, then the operator $M_{\mathcal{B}}$ is of strong type $(p,p)$ for any $p>1$. More precisely, for $i=1,\dotsc,d$ let
$\Omega_{i}:=\{(n_{ki},l_{k,i}):k\in\mathbb{N}\}$
and for $\alpha>0$ define
$\Omega_{i}^{(\alpha)}:=\{(x,y)\in\mathbb{Z}\times\mathbb{N}:|x-n|\leq\alpha(y-l)\mbox{ for some }(n,l)\in\Omega_{i}\}.$
Finally, given $\lambda\in\mathbb{N}$ put
@@
Then the following holds.
Theorem 1.1.
Assume that for all $i=1,\dotsc,d$
($C_{i}$) $\exists\,A,\alpha>0:\quad\#\Omega_{i}^{(\alpha)}(\lambda)\leq A\lambda\quad\mbox{ for all }\lambda\in\mathbb{N}.$
Then the operator ${M_{\mathcal{B}}}$ is of strong type $(p,p)$ for any $p>1$.
@@
It does not follow from our proof that $M_{\mathcal{B}}$ is of weak type $(1,1)$. In particular, it appears that extending the ideas of Bellow, Jones, and Rosenblatt to prove a weak type-$(1,1)$ inequality for $\mathbb{Z}^{d}$- or $\mathbb{R}^{d}$-actions with $d\geq 2$ is a non trivial task. This is mainly due to obstructions in the use of the Hardy-Littlewood Maximal Inequality in higher-dimension (see [BJR90, pages 45 and 46]).
As a corollary to Theorem 1.1 we deduce that, under Condition ($C_{i}$) for $i=1,\dotsc,d$, the sequence of functions $A_{k}f$ converges point-wise for any $f\in\textup{L}^{p}(X)$ with $p>1$.
Corollary 1.3.
Assume that the action generated by the transformations $T_{1},\dotsc,T_{d}$ on the space $X$ is ergodic and that Condition ($C_{i}$) is satisfied for all $i=1,\dotsc,d$. Then for any $f\in\textup{L}^{p}(X)$ with $p>1$ and $\mu$-a.e. $x\in X$ we have that
$A_{k}f(x)\to\mu(f)\quad\mbox{as }k\to\infty.$
If Condition ($C_{i}$) is not satisfied for at least one $1\leq i\leq d$, on the other hand, and the action of $\mathbb{Z}^{d}$ on $X$ is aperiodic, the conclusion of Corollary (1.3) fails in a strong sense. Recall that a measurable action of $\mathbb{Z}^{d}$ on $X$ is said to be aperiodic if $\mu(x\in X:\boldsymbol{a}.x=x)=0$ for all $\boldsymbol{a}\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}$.
Before formally stating the converse of Corollary 1.3, let us also recall the definition of mixing family introduced by Sawyer. Following [Saw66], we say that a family of measure-preserving transformations $\{S_{h}\}$ on $X$ is mixing if for any pair of measurable subsets $A,B$ of $X$ and any $\rho>1$
Remark 1.4.
$\inf_{h}\mu\left(A\cap S_{h}(B)\right)\leq\mu(A)\mu(B).$
Compare this to
$\lim_{S_{h}\to\infty}\mu\left(A\cap S_{h}(B)\right)=\mu(A)\mu(B),$
which corresponds the usual definition of mixing action. In fact, it is possible to show that any group of measure-preserving transformations that acts ergodically on $X$ satisfies (1.1) (see, e.g., [Saw66, Lemma 1, page 163]). This is an easy exercise if $\{S_{h}\}$ is just the group generated by one ergodic transformation on $X$.
We are now in a position to state our next result.
Theorem 1.5.
Assume that the action generated by the transformations $T_{1},\dotsc,T_{d}$ on the space $X$ is aperiodic and that $T_{1},\dotsc,T_{d}$ commute with a mixing family $\{S_{h}\}$ on $X$. Then, if for some $1\leq i\leq d$ it holds that
($!C_{i}$) $\forall\,A,\alpha>0\quad\#\Omega_{i}^{(\alpha)}(\lambda)\geq A\cdot\lambda\quad\mbox{for an unbounded set of }\lambda>0,$
the operators $A_{k}$ have the "strong sweeping out" property, that is, for any $\varepsilon>0$ there exists a set $B_{\varepsilon}\subset X$, with $\mu(B_{\varepsilon})<\varepsilon$, such that for $\mu$-a.e. $x\in X$ it holds that
$\liminf_{k}A_{k}\chi_{B_{\varepsilon}}(x)=0\quad\mbox{and}\quad\limsup_{k}A_{k}\chi_{B_{\varepsilon}}(x)=1.$
Remark 1.6.
It is worth pointing out that, in view of Remark 1.4, any subgroup of an Abelian group acting ergodically on the space $(X,\mu)$ commutes with a mixing family, even if the subgroup itself does not act ergodically.
@@ 1.2. The Continuous Case
We will now state the continuous version of the results of the previous subsection.
@@
and let :=$\mathcal{B}:=\{B_{k}\}$. To any box $B_{k}\in\mathcal{B}$ associate a continuous averaging operator $R_{k}$ defined by
@@
Then, in analogy to Theorem 1.1, we have the following.
@@
Once again, the analog of Corollary 1.3 holds.
@@
We now proceed to state the following converse to Corollary 1.8. Recall that a $d$-dimensional flow $U_{\boldsymbol{t}}$ on $X$ is said to be aperiodic if there exists a set $N\subset X$, with $\mu(N)=0$, such that $U_{\boldsymbol{t}}x\neq x$ for all $x\in X\setminus N$ and all $\boldsymbol{t}\in{}^{d}\setminus\{\boldsymbol{0}\}$.
@@
($!\tilde{C}_{i}$) $\forall\,A,\alpha>0\quad\textup{Leb}\left(\Omega_{i}^{(\alpha)}(\lambda)\right)\geq A\cdot\lambda\quad\mbox{for an unbounded set of }\lambda>0,$
the operators $R_{k}$ have the "strong sweeping out" property (see Theorem 1.5).
In the next subsection, we present an application of Theorem 1.9.
@@ 1.3. Failure of Submanifold Genericity for Essentially Bounded Functions
Let $M\subset\mathbb{R}^{d}$ be a compact $m$-dimensional $\mathscr{C}^{1}$ submanifold of $\mathbb{R}^{d}$ and let $\mathcal{F}\subset\textup{L}^{1}(X)$ be a collection of functions. Let $\textup{vol}_{m}$ denote the $m$-dimensional volume measure induced by the Euclidean metric on $\mathbb{R}^{d}$. We say that a measure $\nu$ on $X$ (potentially equal to $\mu$) is $(M,\mathcal{F})$-generic if for $\nu$-a.e. $x\in X$ it holds that
@@
When $M$ is a $d$-dimensional bounded positive-measure set, $(M,\textup{L}^{1}(X))$-genericity is equivalent to ergodicity of the $\mathbb{R}^{d}$-action (see [Lin01, Theorem 1.2]). However, if $M$ is a proper submanifold of $\mathbb{R}^{d}$, the convergence in (1.2) is more delicate and may not occur for all $f\in\textup{L}^{1}(X)$. For example, when $M=\mathbb{S}^{d-1}$ (i.e., the $(d-1)$-dimensional sphere in $\mathbb{R}^{d}$) and $d\geq 3$, Jones [Jon93, Theorem 2.1] showed that $\mu$ is $(M,\textup{L}^{p}(X))$-generic for any $p>d/(d-1)$ and that the bound on $p$ is sharp (see [Jon93, Theorem 2.3] for a constructive counterexample). [Lac95] extended this result to the case $d=2$. Such analytical arguments, however, do not apply to different manifolds $M$ such as, e.g., the boundary of a hypercube.
If one restricts to very special families of functions $\mathcal{F}$, on the other hand, there are actions for which point-wise convergence in (1.2) occurs for essentially any sufficiently regular submanifold of $\mathbb{R}^{d}$. For example, let $X=G/\Gamma$, where $G$ is a semisimple Lie group (with no factors of rank $1$) and $\Gamma$ is a lattice in $G$. Let $\mu$ denote the canonical left-invariant measure on $G/\Gamma$ and let $A<G$ be a Cartan subgroup, whose Lie algebra will be denoted by $\mathfrak{a}$. Consider the action of $\mathfrak{a}$ on $X$ given by $x\mapsto\exp(a)x$ (which is identifiable with an $\mathbb{R}^{d}$ action on $X$, if $G$ is of rank $d$). Then [BFK25, Theorem 1.3] asserts that the measure $\mu$ on $X$ is $(k,\mathscr{C}^{\infty}_{c}(X))$-generic for all $1\leq k\leq d$. Here, $\mathscr{C}^{\infty}_{c}(X)$ stands for the space of smooth compactly supported functions on $X$. In addition, any invariant measure supported on a closed unipotent orbit in $X$ is also $(k,\mathscr{C}^{\infty}_{c}(X))$-generic for all $k=1,\dotsc,d$ (see for instance [BFK25, Theorem 1.4]). More generally, any measure-preserving and exponentially mixing action of $\mathbb{R}^{d}$ on a probability space $(X,\mu)$ (under some minor assumptions) enjoys the same property (see [BFK25, Theorem 2.2]).
In view of the above discussion, it is natural to ask whether there exist ergodic actions of $\mathbb{R}^{d}$ on some probability space $(X,\mu)$ such that the measure $\mu$ is $(k,\mathcal{F})$-generic for some much larger class of functions $\mathcal{F}$, e.g., $L^{\infty}(X)$. As a consequence of Theorem 1.9, we are able to show the following.
@@
Let $(X,\mu)$ be a probability space, equipped with an ergodic and aperiodic measure-preserving action of $\mathbb{R}^{d}$. Let $M$ be a compact $m$-dimensional $\mathscr{C}^{1}$ submanifold of $\mathbb{R}^{d}$ such that $M\cap\pi$ is a non-empty open set in $M$ for a given $m$-dimensional affine subspace $\pi$ of $\mathbb{R}^{d}$, which does not contain the origin. Then the measure $\mu$ is not $(M,\textup{L}^{\infty}(X))$-generic. In particular, there exist no ergodic and aperiodic measure-preserving actions of $\mathbb{R}^{d}$ on $X$ that are $(m,\textup{L}^{\infty}(X))$-generic for any $m=1,\dotsc,d-1$.
@@
The left-multiplication action of a Cartan subgroup $A$ of a Lie group $G$ on the quotient $X=G/\Gamma$ is always ergodic and aperiodic (see explanation below). Thus, Theorem 1.10 implies that the $G$-invariant measure $\mu$ on $X$ is not $(m,\textup{L}^{\infty}(X))$-generic for any $m=1,\dotsc,d-1$. In other words, a point-wise ergodic theorem for dilates of submanifolds $M$ such as those in Theorem 1.10 cannot hold. However, it is interesting to observe that for the same submanifolds the convergence in (1.2) still occurs in the $\textup{L}^{p}$-norm for any $p\geq 1$, i.e., a mean ergodic theorem holds. This follows easily from the fact that the measure $\mu$ is $(m,\mathscr{C}^{\infty}_{c}(X))$-generic, by a density argument.
Let us briefly explain why the left-multiplication action of a Cartan subgroup $A$ on $G$ is aperiodic. If $ag\Gamma=g\Gamma$ for some $a\in A$ and $g\in G$, then there exists $\gamma\in\Gamma$ such that $g^{-1}ag=\gamma$. Put $S(A,\gamma):=\{g\in G:g\gamma g^{-1}\in A\}$. Hence, any periodic point $g\Gamma$ for the group $A$ belongs to the set
@@
This is a countable union of algebraic varieties and therefore has measure 00.
We conclude with the following observation. If $M$ is a polynomial curve in d not entirely contained in a proper affine subspace, then, according to [BF09, Theorem 2] (with a slight modification to account for the density induced by the measure $\textup{vol}_{m}$), the convergence in (1.2) occurs in norm (more precisely, in the $\textup{L}^{2}$ norm). As suggested in [BF09], it would be natural to study point-wise convergence as in (1.2) for such curves. However, polynomial curves fall outside the scope of Theorem 1.10, since any smooth submanifold $M$ of $\mathbb{R}^{d}$ not entirely contained in, but intersecting an affine subspace in an open set, cannot be analytically parametrized.

### v2

@@ Abstract.
We generalize results of Jones and Olsen on multi-parameter moving ergodic averages to measure-preserving actions of $\mathbb{R}^{d}$ for $d\geq 1$. In particular, we give necessary and sufficient conditions for the pointwise convergence of averages over families of boxes in $\mathbb{R}^{d}$. As an application of our characterization, we show that averages along dilates of "locally flat" submanifolds in $\mathbb{R}^{d}$ do not necessarily converge point-wise for bounded measurable functions. This is closely related to the concept of submanifold-genericity recently introduced in [BFK25].
@@ 1. Introduction
Let $(X,\mu)$ be a non-atomic probability space and let $T_{1},\dotsc,T_{d}$ be commuting measure-preserving transformations on the space $X$. Given sequences $\{\boldsymbol{n}_{k}\}\subset\mathbb{Z}^{d}$ and $\{\boldsymbol{l}_{k}\}\subset\mathbb{N}^{d}$ ($k\in\mathbb{N}$), put
@@
and let $\mathcal{B}:=\{B_{k}\}$. To any box $B_{k}\in\mathcal{B}$ associate an averaging operator $A_{k}$ defined by
@@
In [BJR90], Bellow, Jones, and Rosenblatt gave necessary and sufficient conditions on the sequence $\mathcal{B}$ for $M_{\mathcal{B}}$ to satisfy a maximal inequality in dimension $d=1$. Jones and Olsen [JONES1991127] generalized the result to the case $d\geq 1$.
Let us briefly recall their characterization (see [JONES1991127, Theorem 2.1]). For each $i=1,\dotsc,d$ let
$\Omega_{i}:=\{(n_{ki},l_{k,i}):k\in\mathbb{N}\}.$
For fixed $\alpha>0$ and $\lambda\in\mathbb{N}$ define
$\Omega_{i}^{(\alpha)}:=\left\{(x,y)\in\mathbb{Z}\times\mathbb{N}:|x-n|\leq\alpha(y-l)\mbox{ for some }(n,l)\in\Omega_{i}\right\}$
and
@@
Then, according to [JONES1991127, Theorem 2.1 Part (a)], $M_{\mathcal{B}}$ is of strong type $(p,p)$ for any $p>1$ if and only if the following condition on the sequence $\mathcal{B}$ is satisfied for all coordinates $i=1,\ldots,d$
($C_{1}$) $\exists\,A,\alpha>0:\quad\#\Omega_{i}^{(\alpha)}(\lambda)\leq A\lambda\quad\mbox{for all }\lambda\in\mathbb{N}.$
Condition ($C_{1}$) asserts, in other words, that the horizontal cross-section of all the cones of aperture $\alpha$ above points in the sequence $\Omega_{i}$ at fixed height must grow linearly as the height increases (see picture below). For example, the sequence $(k,rk)$ for any $r\in\mathbb{N}$ has this property, while the sequence $(k,\sqrt{k})$ does not.
[…]
Figure 1. Set $\Omega_{i}^{(1)}(\lambda)$ for some sequence $(n_{ki},l_{ki})$.
@@
Note that when $d=1$, [BJR90][Theorem 1 Part (a)] shows that $M_{\mathcal{B}}$ is of weak type $(1,1)$. For $d\geq 2$, this does not follow from the proof in [JONES1991127]. In fact, as discussed in [JONES1991127, Remark 1], there exist sequences of boxes $\mathcal{B}$ satisfying ($C_{1}$) such that $M_{\mathcal{B}}$ is not of weak type $(1,1)$.
If the operator $M_{\mathcal{B}}$ is of strong type $(p,p)$ and the action generated by the transformations $T_{1},\ldots,T_{d}$ is ergodic, it follows from a standard argument the sequence of functions $A_{k}f$ converges point-wise for any $f\in\mathrm{L}^{p}$. [JONES1991127, Theorem 3.2] shows additionally that this is not the case if Condition ($C_{1}$) fails in at least one coordinate.
In this paper, we generalize [JONES1991127, Theorem 2.1, Part (a)] and [JONES1991127, Theorem 3.2] to measurable actions of $\mathbb{R}^{d}$ for $d\geq 1$. When $d=1$, an alternative proof can be found in [Ferrando_1995][Section 4] as a special case (sequence of moving cubes of dimension 1). Our main motivation for considering such theorems is to study the property of submanifold genericity, recently introduced in [BFK25]. In particular, based on our extension of Jones and Olsen’s result we are able to show that in aperiodic actions of $\mathbb{R}^{d}$ an ergodic theorem for dilates of general submanifolds is not to be expected, unless the test functions are very regular (e.g., smooth). We now proceed to lay out our results in greater detail.
@@ 1.1. The Continuous Case
and let $\mathcal{B}:=\{B_{k}\}$. To any box $B_{k}\in\mathcal{B}$ associate a continuous averaging operator $R_{k}$ defined by
@@
Then, in analogy to [JONES1991127][Theorem 2.1 Part (a)], we have the following.
@@
As a corollary of 1.2, we show the following.
@@
If Condition ($\tilde{C}_{i}$) is not satisfied for at least one $1\leq i\leq d$, on the other hand, and the action of $\mathbb{R}^{d}$ on $X$ is aperiodic, the conclusion of Corollary 1.3 fails in a strong sense (cf. [JONES1991127, Theorem 3.2]). Recall that a $d$-dimensional flow $U_{\boldsymbol{t}}$ on $X$ is said to be aperiodic if there exists a set $N\subset X$, with $\mu(N)=0$, such that $U_{\boldsymbol{t}}x\neq x$ for all $x\in X\setminus N$ and all $\boldsymbol{t}\in\real^{d}\setminus\{\boldsymbol{0}\}$.
Before formally stating the converse of Corollary 1.3, we recall the definition of mixing family introduced by Sawyer. Following [Sawyer], we say that a family of measure-preserving transformations $\{S_{h}\}$ on $X$ is mixing if for any pair of measurable subsets $A,B$ of $X$ and any $\rho>1$
Remark 1.4.
$\inf_{h}\mu\left(A\cap S_{h}(B)\right)\leq\mu(A)\mu(B),$
which differs from the hypothesis that in a mixing action
$\lim_{S_{h}\to\infty}\mu\left(A\cap S_{h}(B)\right)=\mu(A)\mu(B).$
In fact, it is possible to show that any group of measure-preserving transformations that acts ergodically on $X$ satisfies (1.1) (see, e.g., [Sawyer, Lemma 1, page 163]). This is an easy exercise if $\{S_{h}\}$ is just the group generated by one ergodic transformation on $X$.
We are now in a position to state our next result.
@@
($\tilde{C}_{i}$) $\forall\,A,\alpha>0\quad\textup{Leb}\left(\Omega_{i}^{(\alpha)}(\lambda)\right)\geq A\cdot\lambda\quad\mbox{for an unbounded set of }\lambda>0,$
the operators $R_{k}$ have the "strong sweeping out" property, that is, for any $\varepsilon>0$ there exists a set $B_{\varepsilon}\subset X$, with $\mu(B_{\varepsilon})<\varepsilon$, such that for $\mu$-a.e. $x\in X$ it holds that
$\liminf_{k}R_{k}\chi_{B_{\varepsilon}}(x)=0\quad\mbox{and}\quad\limsup_{k}R_{k}\chi_{B_{\varepsilon}}(x)=1.$
Remark 1.6.
It is worth pointing out that, in view of Remark 1.4, any subgroup of an Abelian group that acts ergodically on the space $(X,\mu)$ commutes with a mixing family, even if the subgroup itself does not act ergodically.
In the next subsection, we present the main application of Theorem 1.5.
@@ 1.2. Failure of Submanifold Genericity for Essentially Bounded Functions
Let $M\subset\mathbb{R}^{d}$ be a compact $m$-dimensional $\mathscr{C}^{1}$ submanifold of $\mathbb{R}^{d}$ and let $\mathcal{F}\subset\textup{L}^{1}(X)$ be an arbitrary collection of functions. Let $\textup{vol}_{m}$ denote the $m$-dimensional volume measure induced by the Euclidean metric on $\mathbb{R}^{d}$. We say that a measure $\nu$ on $X$ (potentially equal to $\mu$) is $(M,\mathcal{F})$-generic if for $\nu$-almost every $x\in X$ it holds that
@@
When $M$ is a $d$-dimensional bounded positive-measure set, $(M,\textup{L}^{1}(X))$-genericity is equivalent to ergodicity of the $\mathbb{R}^{d}$-action (see [Lindenstrauss, Theorem 1.2]). However, if $M$ is a proper submanifold of $\mathbb{R}^{d}$, the convergence in (1.2) is more delicate and may not occur for all $f\in\textup{L}^{1}(X)$. For example, when $M=\mathbb{S}^{d-1}$ (i.e., the (d1)(d-1)-dimensional sphere in $\mathbb{R}^{d}$) and $d\geq 3$, Jones [Jones, Theorem 2.1] showed that $\mu$ is $(M,\textup{L}^{p}(X))$-generic for any p>d/(d1)p>d/(d-1) and that the bound on $p$ is sharp (see [Jones, Theorem 2.3] for a constructive counterexample). [Lacey] extended this result to the case $d=2$. Such analytical arguments, however, do not apply to different manifolds $M$, such as (portions of) the boundary of a hypercube.
If one restricts to very special families of functions $\mathcal{F}$, on the other hand, there are actions for which point-wise convergence in (1.2) occurs for virtually any sufficiently regular submanifold of $\mathbb{R}^{d}$. For example, let $X=G/\Gamma$, where $G$ is a semisimple Lie group (with no factors of rank $1$) and $\Gamma$ is a lattice in $G$. Let $\mu$ denote the canonical left-invariant measure on $G/\Gamma$ and let $A<G$ be a Cartan subgroup, whose Lie algebra will be denoted by $\mathfrak{a}$. Consider the action of $\mathfrak{a}$ on $X$ given by $x\mapsto\exp(a)x$ (which is identifiable with an $\mathbb{R}^{d}$ action on $X$, if $G$ is of rank $d$). Then [BFK25, Theorem 1.3] asserts that the measure $\mu$ on $X$ is $(k,\mathscr{C}^{\infty}_{c}(X))$-generic for all $1\leq k\leq d$. Here, $\mathscr{C}^{\infty}_{c}(X)$ stands for the space of smooth compactly supported functions on $X$. In addition, any invariant measure supported on a closed unipotent orbit in $X$ is also $(k,\mathscr{C}^{\infty}_{c}(X))$-generic for all $k=1,\dotsc,d$ (see for instance [BFK25, Theorem 1.4]). More generally, any measure-preserving and exponentially mixing action of $\mathbb{R}^{d}$ on a probability space $(X,\mu)$ (under some minor assumptions) enjoys the same property (see [BFK25, Theorem 2.2]).
In view of the above discussion, it is natural to ask whether there exist ergodic actions of $\mathbb{R}^{d}$ on some probability space $(X,\mu)$ such that the measure $\mu$ is $(k,\mathcal{F})$-generic for some much larger class of functions $\mathcal{F}$, e.g., $\mathrm{L}^{\infty}(X)$. As a consequence of Theorem 1.5, we are able to show the following.
@@
Let $(X,\mu)$ be a probability space, equipped with an ergodic and aperiodic measure-preserving action of $\mathbb{R}^{d}$. Let $M$ be a compact $m$-dimensional ($1\leq m\leq d-1$) $\mathscr{C}^{1}$ submanifold of $\mathbb{R}^{d}$ such that $M\cap\pi$ is a non-empty open set in $M$ for a given $m$-dimensional affine subspace $\pi$ of $\mathbb{R}^{d}$, which does not contain the origin. Then the measure $\mu$ is not $(M,\textup{L}^{\infty}(X))$-generic. In particular, there exist no ergodic and aperiodic measure-preserving actions of $\mathbb{R}^{d}$ on $X$ that are $(m,\textup{L}^{\infty}(X))$-generic for any $m=1,\dotsc,d-1$.
@@
Note that the left-multiplication action of a Cartan subgroup $A$ of a Lie group $G$ on the quotient $X=G/\Gamma$ is always ergodic and aperiodic (see explanation below), so that both [BFK25, Theorem 1.3] and Theorem 1.7 apply simultaneously. In other words, a point-wise ergodic theorem for dilates of submanifolds $M$ such as those in Theorem 1.7 holds for smooth test functions, but fails as soon as the regularity assumption is dropped (e.g., if the test function is just essentially bounded). It is also interesting to observe that for the same submanifolds the convergence in (1.2) still occurs in the $\textup{L}^{p}$-norm for any $p\geq 1$, i.e., a mean ergodic theorem holds. This follows easily from the fact that the measure $\mu$ is $(m,\mathscr{C}^{\infty}_{c}(X))$-generic, by a density argument.
Let us briefly explain why the left-multiplication action of a Cartan subgroup $A$ on $G$ is aperiodic. If $ag\Gamma=g\Gamma$ for some $a\in A$ and $g\in G$, then there exists $\gamma\in\Gamma$ such that $g^{-1}ag=\gamma$. Put $S(A,\gamma):=\{g\in G:g\gamma g^{-1}\in A\}$ and observe that any periodic point $g\Gamma$ for the group $A$ belongs to the set
@@
This is a countable union (since $\Gamma$ is a lattice in $G$) of algebraic varieties and therefore has measure 0.
We conclude with the following observation. According to [BF09, Theorem 2], if $M$ is a polynomial curve in d not entirely contained in a proper affine subspace, the convergence in (1.2) occurs in norm (more precisely, in the $\textup{L}^{2}$ norm). As suggested in [BF09], it would be natural to study the point-wise convergence in (1.2) for such curves. However, polynomial curves fall outside the scope of Theorem 1.7, since any analytic submanifold $M$ of $\mathbb{R}^{d}$ cannot intersect an affine subspace in an open set, unless it is entirely contained in the subspace.